OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

f(x) = 9x - 2 and g(x) = -x + 3. f(g(x) = 9(-x+3) -2

OpenStudy (wmj259):

This kind of problem is called a composite function. Just to let you know :).